# Exploring Imbalanced Annotations for Effective In-Context Learning

### v1

@@ 4.2 Modeling the conditional bias $\bm{\beta}$
  • Flexible: Our method can be easily incorporated into existing selection methods under the class-imbalanced setting in $\mathcal{D}_{c}$.
  • Easy to use: Our method does not require heavy hyperparameter tuning, as it is insensitive to balance set $\mathcal{D}_{b}$ or class-wise weights $\bm{w}$ (see Figure LABEL:figure:5 (a) and (b)).
  • Model-agnostic: Our method requires access only to the model outputs and integrates effortlessly with any LLMs (see Figure LABEL:figure:5 (c) and (d)).
5 Experiments
5.1 Experimental Setup
Datasets and Evaluation For our evaluations, we verify the effectiveness of our method on four benchmark datasets, including Sentiment Classification (Amazon (Keung et al., 2020), Yelp (Zhang et al., 2015)) and Topic Classification (AgNews, Yahoo (Zhang et al., 2015)). Due to limited space, these tasks’ input/output, statistics and split are reported in the Appendix A.4. To simulate the issue of imbalance data, we generate imbalanced datasetss with pre-defined imbalance ratios $\phi$ (e.g., 0, 10, 50, 100) and we evaluate the performance of ICL on another corresponding class-balanced test dataset. We report average accuracy (%) with standard deviation (over 3 runs).
Models and Baselines We utilize various large language models, including Llama3-8B (AI@Meta, 2024), Mistral-7B (Jiang et al., 2023) and OPT-6.7B (Zhang et al., 2022a). We use Bert-base-uncased sentence encoder (Devlin et al., 2019) as the similarity tokenizer. We also conduct experiments with existing selection methods as baselines, including Random (Min et al., 2022), TopK (Liu et al., 2022), DPP (Ye et al., 2023a), VoteK (Su et al., 2023), ConE (Peng et al., 2024), ByCS (Wang et al., 2024). For hyperparameters, we set the size of the balanced validation set as $|\mathcal{D}_{b}|=100$ by default. The details of our experiments are presented in Appendix A.4.
5.2 Main Results
Can our method improve the performance of existing selection methods? Table 1 presents the average test accuracy of in-context learning with different selection methods on four benchmark datasets using OPT-6.7B, under various imbalance ratios. A salient observation is that our method drastically improves the performance of ICL with imbalanced annotated datasets. For example, for the 100 imbalance ratio, our method improves the average accuracy of six existing selection methods from 37.93 to 43.29 – a 5.46 of direct improvement. More importantly, we show that our method can boost performance for a wide range of selection methods, such as TopK and DPP. For example, we observe that, for the 100 imbalance ratio, the test accuracy of the TopK is improved to 4.57 when integrating our methods. Table 1 also presents the results for different demonstration numbers $K$, suggesting that our method can consistently improve the performance of ICL against imbalanced datasets with varying demonstration numbers $K$. The detailed results are reported in Appendix A.6.3.
@@ 5.2 Main Results
How does the balanced validation set affect the robustness of our method? In Figure LABEL:figure:5 (a), we ablate how the size of balanced validation set $\mathcal{D}_{b}$ affects the effectiveness of our method (cf. Eq. 7). The baselines indicates all candidate demonstrations are selected without our method. It’s noteworthy that our method shows robustness to the choice of the size of the balanced validation set $|\mathcal{D}_{b}|$. Even when we set $|\mathcal{D}_{b}|=50$, it still yields significant improvements in ICL performance on four classification tasks across six selection methods against imbalanced datasets. Due to space constraints, we only report the average results of multiple baselines on various downstream tasks.
@@
Is our method effective with different class-wise weights $\bm{w}$? In the paper, we employ effective numbers as class-wise weights $\bm{w}$, which have been confirmed to empirically successfully deal with imbalance problems in previous studies (Cui et al., 2019; Jamal et al., 2020). We also verify whether the type of class-wise weights affects the performance of our methods. Specifically, we employ class frequency $w_{j}=\frac{n_{j}}{N}$ (Kang et al., 2020; Shi et al., 2023) as an alternative class-wise weights to verify the effectiveness of our method. Our results in Figure LABEL:figure:5 (b) show that our method can improve the ICL’s performance across various imbalanced ratios and achieve similar accuracy to the effective numbers.
Is our method effective with different model architectures? To demonstrate that our proposed method is model-agnostic, we conduct experiments on diverse open-sourced LLMs (Llama3-8B (AI@Meta, 2024) and Mistral-7B (Jiang et al., 2023)) and present the average results of four common benchmarks using the two LLMs in Figure LABEL:figure:5 (c). From the results, we observe that our method consistently improves test performance on various downstream tasks when using different model architectures. For instance, using our method boosts the test accuracy from 47.10 to 54.31, a 7.21 of direct improvement with a 100 imbalance ratio. Therefore, our method can improve the performance of ICL using various open-source LLMs against imbalanced annotated datasets. More details are presented in Appendix A.6.4.
Is our method effective with closed-sourced LLMs? While our method has demonstrated strong promise in open-source LLMs, we also verify its effectiveness using closed-source models. Specifically, we tested our method on ChatGPT-3.5-Turbo across four selection methods (Random, TopK, DPP and VoteK) with various imbalance ratios. Figure LABEL:figure:5 (d) reports the average results for ChatGPT-3.5-Turbo, showing that imbalanced annotated datasets significantly hurt the ICL performance of closed-source LLMs when employing these existing selection methods. After integrating our method, Figure LABEL:figure:5 (d) also shows significant improvements in average test accuracy across four downstream tasks, indicating the effectiveness of our method in addressing imbalances in closed-source LLMs.
@@
Can our method improve real-world imbalanced dataset We verify the effectiveness of our method on a real-world imbalanced dataset. The Emotion (Saravia et al., 2018) dataset has a long-tailed distribution (refer to Figure 1). Figure 18 shown that our method consistently improves existing selection methods and outperforms existing rebalance methods, including over-sampling (Chawla et al., 2002), under-sampling (Liu et al., 2008), stratified sampling (Vilariño et al., 2005), and re-weighting (Cui et al., 2019). For example, with OPT-6.7B (Zhang et al., 2022a), using our method boosts the average accuracy of six selection methods from 21.22 to 23.36, a relative improvement of 10.08%. These results verify that our method is effective in improving ICL’s performance in real-world imbalanced scenarios.
@@ 6.1 Data imbalance in text generation tasks
Table 2 demonstrates that imbalanced annotated dataset significantly hurt ICL’s performance on generation tasks. Additionally, Table 2 also confirms the ineffectiveness of classical data balancing methods, including over-sampling (Chawla et al., 2002), under-sampling (Liu et al., 2008), stratified sampling (Vilariño et al., 2005), and re-weighting (Cui et al., 2019). After integrating our method, the existing selection methods show significant improvements in inference performance, indicating the generalization of our method to text generation. For example, our method improves the average exact match score from 21.20 to 22.93, a notable improvement of 0.73 over six existing demonstration methods. The detailed results of CodeSearchNet (Husain et al., 2019) task are presented in Table 7.
Table 2: The average EM score of NQ dataset across six selection methods with various imbalance ratios. Bold numbers are the best results.
Imbalance Ratio 0 10 50 100
Baselines 23.67 22.47 21.33 21.20
+Over-Sampling 23.60 22.53 22.53 22.33
+Under-Sampling 23.60 19.87 18.47 17.80
+Stratified-Sampling 22.15 20.60 18.82 18.47
+Re-Weighting 23.33 22.60 20.47 20.20
+Ours 24.00 23.87 23.00 22.93
@@ 6.2 Does a larger dataset address the imbalance issue?
Table 3: Average classification accuracy (%) of Baselines / +Ours methods on the Yahoo dataset with various imbalance ratios. Bold numbers are superior results.
Imbalance Ratio 10 50 100
Baseline/+Ours
Random 37.87/38.07 36.20/39.87 33.13/37.73
TopK 47.33/48.87 45.00/47.67 43.07/47.87
DPP 49.60/50.27 48.00/49.33 44.47/47.80
VoteK 47.13/47.80 43.87/47.47 44.67/47.73
ConE 47.53/49.13 45.27/46.80 41.47/45.53
ByDC 48.40/49.33 44.27/47.47 42.17/46.50
Average 46.64/47.25 43.77/46.44 41.50/45.53
While our analysis has shown that imbalanced annotated datasets negatively impact in-context learning by reducing the number of examples in tail classes, one might wonder if this negative effect is due to the reduced size of the dataset rather than a shift in class prior distribution. To address this, we verify the negative impact of imbalanced annotated datasets and the effectiveness of the proposed method on such datasets with an increasing number of examples. Specifically, we consider creating an imbalanced annotated dataset by increasing the number of examples in head classes while keeping the number of examples in tail classes constant. For example, in the case of Yahoo (Zhang et al., 2015), when the imbalance ratio $\phi=0$, the number of annotated examples for each class remains 50. When the imbalance ratio $\phi=100$, the number of examples in the head classes is 5,000, but the number in the tail classes remains 50. We report the test accuracy of Yahoo using OPT-6.7B with six existing selection methods.
@@
7 Conclusion
In this paper, we study the effect of the imbalanced annotated dataset on the performance of in-context learning (ICL). Both our theoretical analysis and empirical investigations reveal that imbalanced annotated datasets significantly hurt the performance of ICL. Meanwhile classical rebalance methods cannot effectively mitigate the imbalanced datasets. To address the issue, we introduce our method, a general strategy that can universally enhance the performance of ICL across various imbalanced ratios. Our method is motivated by decomposing the distributional differences between annotated and test datasets into class-wise weights and conditional bias. Extensive experiments demonstrate that our method can improve the performance of existing selection methods of ICL across various imbalance ratios. Our approach is practical to use, as it is insensitive to the hyperparameters and integrates effortlessly with any LLMs.
Limitations. Our methods need to split a balanced validation dataset from imbalanced annotated datasets, and these methods require that the size of examples in tail classes cannot be too small. Furthermore, it might be challenging to estimate multiple parameters for the class-conditional biases using balanced datasets with limited size. It might be an interesting direction to explore how to model the bias in a more data-efficient manner.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Related Work
@@
In-context learning In-context learning (ICL) is a new paradigm for large language (LLMs), which allows LLMs to make predictions only based on a few demonstrations without explicitly updating parameters (Akyurek et al., 2023; Hendel et al., 2023; Agarwal et al., 2024; Dong et al., 2024; Edwards & Camacho-Collados, 2024; Falck et al., 2024). Many studies show that ICL can achieve performance similar to fine-tuning but without the high computational cost (Gonen et al., 2023; Mosbach et al., 2023; Müller et al., 2024; Panwar et al., 2024). Despite achieving such outstanding performance, ICL has been criticized for being very sensitive to the quality of in-context examples (Fei et al., 2023; Gao et al., 2024a). Various approaches have been proposed to improve the robustness of ICL in recent years, including meta-tuning LLMs (Brunet et al., 2023), calibration (Abbas et al., 2024), demonstration selection (Zhang et al., 2022b; Nguyen & Wong, 2023; Qin et al., 2024; Ye et al., 2023b; Gao et al., 2024b; Luo et al., 2024; Mo et al., 2024), ordering (Lu et al., 2022; Liu et al., 2024), number (Zhang et al., 2025) and formation (Voronov et al., 2024; Yao et al., 2024).
Notably, existing studies often assume that usually assume that annotated dataset and testing data are both i.i.d. (independent and identically distributed) sampled from the same data distribution (Luo et al., 2024; Van et al., 2024). However, this is not always true for real-world scenarios. Specifically, annotated dataset typically exhibit a long-tailed class distribution, where a small portion of classes have a massive number of sample points but the others are associated with only a few examples (Jamal et al., 2020; Schultheis et al., 2024). It is still mysterious how demonstrations sampled from long-tailed distribution affect the performance of ICL on both text classification and generation tasks. In this paper, we show that annotation quality is crucial for ICL in both text classification and generation, where imbalanced datasetss significantly hurt the performance. Different from previous study (Hong et al., 2024), we propose a simple and effective method to enhance the performance of ICL from the perspective of demonstration selection.
Learning with imbalanced datasets Imbalanced dataset is common in many real-world datasets (Cui et al., 2018, 2019; Jamal et al., 2020; Schultheis et al., 2024). The existing approaches to learning with imbalanced datasets can be classified into two types: (1) training imbalance-robust models with imbalance training datasets: designing imbalance-robust loss function (Jamal et al., 2020; Tan et al., 2020; Park et al., 2023; Bhat et al., 2023; Garcin et al., 2022) or designing imbalance-robust model architectures (Long et al., 2022; Pan et al., 2024) to mitigate the issue of imbalanced datasets. However, this method is not suitable for ICL, which usually hypothesizes that users are unable to apply fine-tuning techniques. (2) handling imbalance examples: handling imbalanced datasets is crucial for ensuring balanced performance across all classes. One simple and intuitive approach to deal with the class-imbalanced problem is re-sampling. Under-sampling methods remove examples from the majority classes, which is infeasible under data imbalanced settings (Liu et al., 2008; Wei et al., 2022). The over-sampling method adds repeated examples for the minority classes, usually causing over-fitting to the minority classes (Chawla et al., 2002; Shi et al., 2023). It is noteworthy that these existing methods mainly focus on the fine-tuning setting, and the literature on how to mitigate the effects of imbalanced datasets in in-context learning is limited. This motivated us to design a method for addressing the issue of imbalanced datasets under the setting of ICL.
A.2 Proof
@@
Here, we decompose the term $\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{c}(\mathrm{\mathbf{y}})}$ into two parts: the class-wise weights $\bm{w}=$, and the conditional weight $\bm{\widetilde{\beta}}=$. The class-wise weights measure the difference between class prior distributions $P(Y)$, and the conditional weight measures the difference between conditional distributions $P(X|Y)$. Then we have
@@
where $\bm{w}$ denotes class-wise weights and conditional weight $\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1$ measures the difference of conditional distribution between the annotated and test datasets.
@@
$\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right](\bm{w}+\bm{\beta}).$ (10)
where conditional bias $\bm{\beta}=\bm{w}\times\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{y}})}\left(\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1\right)$ denotes the deviation of class-wise weights caused by the presence of conditional weight $\bm{\widetilde{\beta}}$.
A.3 Bayesian Optimizing Conditional Bias
In this section, we will introduce how to use Bayesian optimization (Gardner et al., 2014; Nogueira, 2014) to find the optimal value of conditional bias $\beta$ by maximizing empirical $\operatorname{Error}$ as follows:
$\mathcal{F}_{\theta}\left(\beta|\mathcal{D}_{r},\mathcal{D}_{b},\bm{w}\right)=\frac{1}{|\mathcal{D}_{b}|}\sum_{i=1}^{|\mathcal{D}_{b}|}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{(w_{i}+\beta)\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{b})\right\}_{i=1}^{|\mathcal{D}_{r}|}\right),\mathrm{\mathbf{x}}_{b}\right),\mathrm{\mathbf{y}}_{b}\right].$ (11)

### v2

@@ 5.2 Main Results
How does the balanced subset affect the robustness of our method? In Figure 3 (a), we examine how the size of the balanced subset $\mathcal{D}_{b}$ affects the effectiveness of our method (cf. Eq. 9). The Vanilla indicates all candidate demonstrations are selected without our method. It’s noteworthy that our method shows robustness to the choice of the size of the balanced subset $|\mathcal{D}_{b}|$. Even when we set $|\mathcal{D}_{b}|=50$, it still yields significant improvements in ICL performance on AgNews and Yahoo datasets across six selection methods against imbalanced datasets. Notably, the balanced subset can be reused in the process of demonstration selection. Thus, our method is applicable in scenarios with extremely limited data from tail classes.
@@
Is our method robust with different class weights $\bm{w}$? In the paper, we employ effective numbers as class weights $\bm{w}$, which have been confirmed to successfully deal with imbalance problems in previous studies [9, 28]. We employ class frequency $w_{j}=\frac{n_{j}}{N}$ [24, 32] as an alternative class weights to verify whether the type of class weights affects the performance of our methods. Figure 3 (b) shows that our method with class frequency can improve the ICL’s performance across various imbalance ratios and achieve accuracy similar to that of effective numbers.
Does our method work with the different number selected candidates $K^{\prime}$? We set a large value of $K^{\prime}$ to ensure that all classes (especially tailed classes) are included in the subset of candidates selected by existing methods. Given a small $K^{\prime}$, the candidate subset may contain too few examples from tailed classes (even missing), leading to suboptimal performance. Therefore, the $K^{\prime}$ should be larger in cases of larger imbalanced ratios. We validate the claim by conducting experiments with varying $K^{\prime}$ on datasets with various imbalanced ratios. The result in Figure 3 (c) demonstrates that one should set a sufficiently large $K^{\prime}$ to achieve the best performance.
Is our method effective with different model architectures and sizes? To show our proposed method is model-agnostic, we conduct experiments on a diverse collection of model architectures and sizes, including open-weight models: OPT-6.7B, OPT-13B, OPT-30B, LLAMA-3-8B and LLAMA-3-70B; and APIs: ChatGPT-3.5-Turbo and Gemini-2.0-Flash. The results (vanilla/+ours) in the Figure 3 (d) present the same phenomenon as the main experiments in the manuscript: the ICL performance of LLMs gets worse at larger imbalanced ratios, and our method can significantly improve the performance, especially at large imbalanced ratios (e.g., 100). For instance, for LLAMA-3-8B [20], using our method boosts the average accuracy from 65.72 to 73.20, a 7.48 of direct improvement with a 100 imbalance ratio. The detailed results of each model are presented in the Table 2.
@@
Can our method work in real-world imbalanced datasets? We verify the effectiveness of our method on a real-world imbalanced dataset. In Emotion [12], a few sentiments make a large contribution and data tend to show a long-tailed distribution (refer to Figure 5).. For example, “joy“ and “sadness“ are head classes, while most other classes, such as “love“ and “surprise“, are tail classes. Table 3 shows that our method consistently improves existing selection methods and outperforms existing rebalancing methods. For example, with OPT-6.7B [19], using our method boosts the average Macro-F1 of six selection methods from 32.12 to 36.69, a direct improvement of 4.57. These results verify that our method is effective in improving ICL’s performance in real-world imbalanced scenarios.
Figure 3: Results for ablation studies. (a) Ablation on the different balanced subset sizes. (b) Ablation on the different class weights (c) Ablation on the different selected candidates. (d) Ablation on the different model architectures and sizes.
@@ 6.1 Data imbalance in text generation tasks
[…]
Figure 4: The impact of imbalanced annotations on the performance of In-Context Learning (ICL) in generation tasks, including Natural Questions (NQ) (a) and CodeSearchNet (b). Figure (c) shows the result of imbalanced datasets with an increasing number of examples on four classification tasks.
@@ 6.2 Does a larger dataset address the imbalance issue?
While our analysis has shown that imbalanced annotations negatively impact ICL by reducing the number of examples in tail classes, one might wonder if this negative effect is due to the reduced size of the dataset rather than a shift in class prior distribution. To address this, we verify the negative impact of imbalanced annotations and the effectiveness of the proposed method on such datasets with an increasing number of examples. Specifically, we consider creating an imbalanced dataset by increasing the number of examples in the head classes while keeping the number of examples in the tail classes constant. For example, in the case of Yahoo [11], when the imbalance ratio $\phi=1$, the number of annotated examples for each class remains 50. When the imbalance ratio $\phi=100$, the number of examples in the head classes is 5,000, but the number in the tail classes remains 50. We report the test accuracy of four classification datasets with six existing selection methods.
Figure 4 (c) shows that imbalanced annotations significantly hurt ICL’s performance when more examples are added to the head classes while keeping the number of examples in the tail classes constant. This illustrates that the decreasing trend in ICL’s performance mainly depends on the class prior distribution of the annotated dataset. Notably, our method significantly improved the ICL’s performance. For instance, for a 100 imbalance ratio, our method outperforms vanilla by a large margin of 5.43, indicating the effectiveness of our method against imbalanced annotation.
7 Conclusion
In this paper, we presented a new research direction for understanding and mitigating the effect of imbalanced annotations on in-context learning (ICL). Our study reveals that imbalanced annotations significantly degrade the performance of ICL across various tasks, while classical rebalancing methods yield limited improvements. We propose RCB, a complementary strategy that estimates conditional bias for reweighting the demonstration score in selection, universally enhancing the performance of ICL with imbalanced annotations. To the best of our knowledge, our method is the first to tackle the class-imbalanced issue from the perspective of demonstration selection. Extensive experiments demonstrate that RCB can enhance the performance of existing selection methods in ICL under class-imbalanced settings. Our approach is easy to use in practice, as it is insensitive to the hyperparameters and does not introduce heavy computational cost.
Limitations. Our methods need to select a balanced subset from imbalanced datasets, which might be impractical if there are few examples in the tail classes. It might be an interesting direction to explore how to model the bias in a more data-efficient manner.
Appendix
@@ 7 Conclusion
Appendix A Related Work
@@
In-context learning In-context learning (ICL) is a new paradigm for large language (LLMs), which allows LLMs to make predictions only based on a few demonstrations without explicitly updating parameters [33, 34, 35, 36, 37, 38]. Many studies show that ICL can achieve performance similar to fine-tuning but without the high computational cost [39, 2, 40, 3]. Despite achieving such outstanding performance, ICL has been criticized for being very sensitive to the quality of in-context examples [41, 42]. Various approaches have been proposed to improve the robustness of ICL in recent years, including meta-tuning LLMs [43], calibration [44], demonstration selection [45, 46, 47, 23, 48, 49, 50], ordering [51, 52], number [53] and formation [54, 55].
The ICL is closely related to the concept of label bias [1], where language models are biased toward certain answers during few-shot learning. In particular, the most relevant part is the majority label bias, a type of context label bias that leads the model to predict answers that appear frequently in the prompt [56, 57]. An empirical work [6] suggested that label imbalance does not substantially impact ICL performance in binary classification. The subsequent works [41, 58, 59, 60] delved deeper into this challenge and proposed to address it by calibrating the model’s output probabilities to compensate for this bias. While prior studies have typically focused on imbalanced distributions within prompts, our work shifts the emphasis to class imbalance in the annotation datasets from which demonstrations are drawn. Notably, this study represents the first systematic attempt to tackle the class-imbalanced issue from the perspective of demonstration selection.
Learning with imbalanced datasets Imbalanced datasets are common in many real-world datasets, so the challenge of class imbalance has been widely studied in the literature [61, 9, 28, 62]. There are two popular directions in learning with imbalanced datasets: (1) Re-sampling: Re-sampling is a widely used strategy in class-imbalanced learning [24], such as Over-sampling [25], which involves repeating data from the minority classes; Under-sampling [26], which involves removing a proportion of data from the majority classes. Stratified sampling [27] samples from each class have an identical probability of being sampled. In this work, we demonstrate that existing rebalancing methods yield limited improvement in in-context learning (ICL) when dealing with imbalanced annotations. (2) Training imbalance-robust models: designing loss function [28, 63, 64, 65, 66] or designing model architectures [67, 68] to mitigate the issue of imbalanced datasets. However, this method inevitably incurs high training costs that might be impractical for LLMs. In contrast, our approach only estimates the conditional bias for each class, resulting in negligible computational costs compared to traditional training methods.
Appendix B Proof
@@
Here, we decompose the term $\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})P_{c}(\mathrm{\mathbf{y}})}$ into two parts: the class weights $\bm{w}=P_{t}(\mathrm{\mathbf{y}})/P_{c}(\mathrm{\mathbf{y}})$, and the conditional weight $\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}$ The class weights measure the difference between class prior distributions $P(Y)$, and the conditional weight measures the difference between conditional distributions $P(X|Y)$. Then we have
@@
where $\bm{w}$ denotes class weights and conditional weight $\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1$ measures the difference of conditional distribution between the annotated and test datasets.
@@
$\mathbb{E}_{P_{c}(\mathrm{\mathbf{x}},\mathrm{\mathbf{y}})}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\left\{s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{t})\right\}_{i=1}^{N}\right),\mathrm{\mathbf{x}}_{t}\right),\mathrm{\mathbf{y}}_{t}\right](\bm{w}+\bm{\beta}).$
where conditional bias $\bm{\beta}=\bm{w}\times\bm{\widetilde{\beta}}=\frac{P_{t}(\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{y}})}\left(\frac{P_{t}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}{P_{c}(\mathrm{\mathbf{x}}|\mathrm{\mathbf{y}})}-1\right)$ denotes the deviation of class weights caused by the presence of conditional weight $\bm{\widetilde{\beta}}$.
Appendix C Estimating Conditional Bias by Bayesian optimization
In this section, we will introduce how to use Bayesian optimization [69, 70] to estimate conditional bias $\bm{\beta}$ as follows:
$\bm{\beta}=\mathop{\arg\min}_{\bm{\beta}}\frac{1}{|\mathcal{D}_{b}|}\sum_{i=1}^{|\mathcal{D}_{b}|}M\left[f_{\theta}\left(\operatorname{Top}_{K}\left(\{(\bm{w}+\bm{\beta})\times s(\mathrm{\mathbf{c}}_{i},\mathrm{\mathbf{x}}_{b})\right\}_{i=1}^{|\mathcal{D}_{r}|}),\mathrm{\mathbf{x}}_{b}\right),\mathrm{\mathbf{y}}_{b}\right],$ (11)